OpenStudy (anonymous):

I need help with this problem √(63x^15y^9/7xy^11)

OpenStudy (anonymous):

since it is a square root,just square root all of the numbers and divide all of the exponents of the variables by 2 for example x^15 will become x^15/2 and x will become x^1/2

OpenStudy (anonymous):

is 3x^7/y the right answer?

OpenStudy (saifoo.khan):

\[\Large \sqrt{ \frac{9x^{14}}{y^2}}\to \frac{3x^7}{y}\]